OpenStudy (anonymous):

write an equation in slope intercept form for a line that passes through the origin with a slope of 3/2

OpenStudy (anonymous):

3y=2x 2y=3x y=3/2x 3x-2y=0

OpenStudy (anonymous):

y = (3/2)x, or y = 1.5 x. The slope-intercept form is y = mx + b, with m as the slope and b as the y-intercept. Sine the origin is (0,0), the y-intercept here is 0. You're given that the slope is 3/2, so there you go :)

OpenStudy (anonymous):

thnks :)

OpenStudy (anonymous):

Answers B and C, when simplified, are the same thing. for 2y = 3x, if you divide each side by 2, you get y = 3/2 x, the same as answer C. and no problem :)